# CardioLive: Empowering Video Streaming with Online Cardiac Monitoring

### v1

@@ Abstract
Online Cardiac Monitoring (OCM) emerges as a compelling enhancement for the next-generation video streaming platforms. It enables various applications including remote health, online affective computing, and deepfake detection. Yet the physiological information encapsulated in the video streams has been long neglected. In this paper, we present the design and implementation of CardioLive, the first online cardiac monitoring system in video streaming platforms. We leverage the naturally co-existed video and audio streams and devise CardioNet, the first audio-visual network to learn the cardiac series. It incorporates multiple unique designs to extract temporal and spectral features, ensuring robust performance under realistic video streaming conditions. To enable the Service-On-Demand online cardiac monitoring, we implement CardioLive as a plug-and-play middleware service and develop systematic solutions to practical issues including changing FPS and unsynchronized streams. Extensive experiments have been done to demonstrate the effectiveness of our system. We achieve a Mean Square Error (MAE) of 1.79 BPM error, outperforming the video-only and audio-only solutions by 69.2% and 81.2%, respectively. Our CardioLive service achieves average throughputs of 115.97 and 98.16 FPS when implemented in Zoom and YouTube. We believe our work opens up new applications for video stream systems. We will release the code soon.
@@
1 Introduction
Video streaming has exploded in recent years, and its growth shows no signs of slowing down. From social platforms like TikTok that have turned live video sharing into a global phenomenon, to Zoom, which has become synonymous with remote work and learning, video streaming has woven itself into the fabric of our daily lives. The popularity of these platforms has not waned even after the COVID-19 pandemic. The market is booming steadily [13], reflecting our collective appetite for real-time, interactive, and accessible content.
Online Cardiac Monitoring (OCM) can be one intriguing enhancement for the next-generation video streaming platforms. The rich tapestry of video and audio in streaming not only provides the context of actions, movement, human activities, speech, etc., but it also embeds subtle cardiac events, which have been long neglected in contemporary video streaming systems. Uncovering such physiological information would bring various benefits. In the realm of remote health, physicians could remotely access real-time cardiac data without the need for specialized equipment [4]. Similarly, in video gaming, displaying a player’s heart rate during live streams could add a new layer of excitement and engagement for viewers [9]. Notably, in the Paris Olympics 2024, NBC introduced heart-rate streaming to add a new ”gamifying” element for creating compelling TV [7]. This technology could also play a pivotal role in online conferences or interviews, where emotional responses (including lies) inferred from cardiac data [57, 53] could enrich interactions, making them more nuanced and meaningful. Furthermore, the potential for this technology extends into security and fraud detection against digital impersonation techniques like deepfakes [38, 47]. These multifaceted applications of OCM underscore its potential to revolutionize video streaming, making it not just a tool for communication and entertainment, but also a platform for health monitoring, affective computing, emotional intelligence, and security.
However, existing online cardiac monitoring either relies on specified hardware [8](e.g., heartbeat belt) or introduces additional sensing modalities (e.g., Wi-Fi [39], mmWave [70], and UWB [22] etc.) which are not typically available in live streaming systems. These approaches suffer from extra cost and are often misaligned with live streams. Moreover, sensing-based approaches necessitate active transmission of the sensing signals [58, 48], which is often impractical to force in live streaming applications. A video streaming system that seamlessly enables online cardiac monitoring in pervasive contexts without additional hardware still lacks.
@@ 1 Introduction
First, what information should we take from the video streaming system to monitor the cardiac activities? Existing works [21, 40, 45, 34, 72, 73, 74, 41, 77] on extracting heart rate from human faces focus on remote photoplethysmography (rPPG) which leverages solely video. These video-only solutions are more likely to suffer from low illumination conditions, head movement, and orientation. Recent progress in cardiac Vocal User Interfaces (VUIs) [67] inspires us to infer heart rate from human speech. However, audio signals are usually sensitive to noise interference and lack contextual background information, rendering them less robust in real-life scenarios and requiring user calibration. Conceptually, video provides detailed visual context while sound exhibits resilience to varying light conditions and body motions. Consequently, they offer complementary advantages to enhance cardiac monitoring. This motivates us to move beyond video-only or audio-only solutions, and investigate new designs to combine the naturally co-existed video and audio streams.
@@
Second, how to tackle real-world problems to make this system robust and accurate? Unveiling the cardiac activity from video and audio is challenging. The information is nuanced and easy to be overshadowed by more prominent body movements, environmental dynamics, and/or ambient noise. Previous works [40, 45, 34, 74, 41, 77] primarily evaluate models on well-controlled datasets featuring static subjects under optimized light conditions and viewing angles, which simplifies the problems yet becomes unrealistic in real-world settings. The task gets even more challenging when deployed in live video streaming environments, due to the discrepancies in frame rates, degraded image quality, and presence of multiple individuals with mixed audio and video streams. To deliver an accurate and robust system in practice, novel techniques are desired to effectively discern subtle cardiac signals amidst various disturbances while combating fluctuating frame rates and drifted misalignment of the streams.
@@
Third, how to enable Service-On-Demand (SoD) cardiac monitoring in video streaming systems? Despite the promise of the integration, enabling SoD for users poses significant challenges due to the complexity of modern video streaming systems. These platforms vary widely, encompassing formats such as conferences [10, 12, 6], Video-On-Demand (VoD) [11, 5], live streaming [14, 2], etc, each with its own technical and operational nuances. These providers must balance the demands of real-time data processing with the need for immediate accessibility and minimal latency while not interfering with the original streams. At the same time, deploying our service on edge (e.g., browsers) benefits from preserving privacy, while getting access to the data yields another challenge. One naive way is to deploy our models over the WebRTC peers, but it lacks scalability and versatility. To this end, we are motivated to establish a plug-and-play service that can be seamlessly integrated into video streaming systems, whether hosted on servers or edges.
In this paper, we present CardioLive, the first-of-its-kind online cardiac monitoring system, that can continuously infer the heart rate in video streaming systems. At the core of CardioLive, we design a novel audio-video deep learning network, CardioNet, that can effectively learn the nuanced cardiac activities from facial regions and human voices. Specifically, we combine the temporal difference network and a frequency-aware block to model the temporal-spectrum properties from videos. We directly exploit the raw audio to capture the cardiac activities by emulating the natural filtering effects of the human body. To handle the irregularly sampled data, we integrate time embeddings to provide temporal context. Finally, we fuse audio and visual features through a multi-head temporal attention mechanism, which synergistically combines the strengths of both modalities to produce a robust and precise cardiac monitoring solution.
@@
We further devise systematic solutions to deploy CardioLive as a middleware service to support the SoD online cardiac monitoring. We introduce practical techniques to handle issues like changing FPS and unsynchronized streams. Through in-depth analyses of mainstream video streaming architectures, we realize a CardioLive service with effective data hooks and novel packet and buffer designs, which can be easily integrated with various video streaming systems.
@@
Extensive experiments have been done to validate the effectiveness of CardioLive. We have self-collected data through 8 different devices and 10 users. Our evaluation results show that CardioLive achieves a mean absolute error (MAE) of 1.79 BPM and root mean square error (RMSE) of 3.25 BPM, largely outperforming the video-only solutions by 69.2% in MAE and 61.4% in RMSE, and the audio-only solution by 81.2% in MAE and 76.8% in RMSE. We demonstrate CardioLive’s generalizability to different environments, devices, and users. As for CardioLive service, we implement our system on two ends, a meeting platform (Zoom) and a content provider (YouTube), respectively. We achieve the overall throughput of 115.97 FPS and 98.16 FPS for each platform respectively, ensuring smooth updates without disrupting the original streams. These results highlight the robustness and accuracy of CardioLive, confirming its potential for widespread application in video streaming systems.
@@
❸ We implement CardioLive as a service-based plug-and-play middleware, that can seamlessly be integrated into mainstream platforms for real-time streaming.
2 Design Scope
@@
  • Accessibility: In many video streaming scenarios, such as live product demonstrations on TikTok or Zoom interviews, using wearables or additional hardware is often impractical. OCM can overcome this problem by leveraging modalities that already exist within video streams, thereby increasing accessibility for audiences and facilitating broader engagement. It also promises wider dissemination of remote health, offering device-free cardiac monitoring compared to the latest work [20] that relies on earphones.
@@
  • Enhanced Analytical Abilities: While there exist alternative approaches for tasks including affective computing [44, 46, 66, 16] and deepfake detection [69, 24], the cardiac signal shows a strong correlation with them [63, 46], by capturing the subtle changes in heart rate. In this context, OCM provides an additional verification layer in a real-time and continuous manner, allowing experts to proceed to analyze behaviors. This analysis can help determine if someone is lying, happy, nervous, or engaging in deceptive behavior.
@@
  • Entertainment: Our work also presents a distinct chance for augmented entertainment. With the rise of live streaming, the audience can access the heart rates of celebrities, which opens up a new world for the existing viewing experiences.
@@
Despite the potential, there are no existing solutions capable of achieving this integration without additional hardware. In this work, we focus on addressing this gap by leveraging the co-existence of audio and video signals, specifically in scenarios where a speaker is talking. This can be common in both entertainment and telehealth use cases, including affective computing, remote health, deepfake detection, etc. At the core of OCM is the accurate prediction of cardiac information. Our system should robustly detect the heart rate from the video streaming systems by hooking the video and audio chunks. Once cardiac data is acquired, it can be further analyzed for various downstream tasks, including affective computing, remote health monitoring, and deepfake detection. Yet how cardiac monitoring is used for downstream tasks (e.g., emotions, lies, etc) is not the focus of this paper.
Audio-Video Pair: We intend to integrate the video and audio information for cardiac monitoring. Leveraging the natural co-existence of audio and video modalities offers contemporary benefits as follows:
  • Ubiquity: Video and audio streams are the most fundamental components in video streaming systems, while no additional hardware is needed.
  • Feasibility: Both video and audio data contain the cardiac information (discussed in §3.1).
  • Complementarity: Audio and video offer different strengths and weaknesses. Audio is less interfered with by motion and light but is sensitive to noises. Video is more robust to noises but will fail in various body movements and non-optimized view angles. We will elaborate the detailed analyses in §3.
We argue that in our primary target application scenarios—such as video conferencing, live streaming, and remote healthcare—human speech is inherently present alongside video. Our goal is to fully leverage the potential of these naturally coexisting signals. Additionally, our system is welldesigned to seamlessly fall back to a video-only solution when audio quality degrades.
@@
CardioLive as a service: To deploy such an OCM system, a straightforward way is to build a self-hosted WebRTC service, which, however, does not scale to existing video streaming systems. The recent rise of Software as a service (SaaS) provides a scalable and robust software design paradigm to construct our system. Therefore, for the sake of versatility, we aim to establish a microservice to host CardioLive for seamless integration with mainstream video streaming platforms.
@@
In a nutshell, the audio-video pair appears to be an attractive choice for ubiquitous and practical OCM, yet it entails numerous challenges to build an accurate and robust multi-modal algorithm and system. We will present our model design in §3 and leave the system implementation in §4.
@@
Figure 3: The performances of video-based approaches vary under different body movements light conditions.
3 CardioNet Design
In this section, we will present our design of CardioNet. We will first describe the underlying fundamentals of inferring cardiac activity from video and audio. Then, we will illustrate our design of model.
3.1 Kinetics for Cardiac Learning
Principles: In this section, we introduce the principles of extracting cardiac information from video and audio data. The fundamental concept revolves around the variations in blood pressure caused by cardiac activities, which manifest as quasi-periodic deformations of blood vessels. Since blood vessels circulate blood throughout the body, including the face, lungs, and throat, we can infer heart activity in these areas through video and audio analysis. Specifically, when a light illuminates the skin, subtle color variations caused by pulse-induced blood flow can be captured through video streams. Additionally, as the lungs supply airflow for vocal fold vibrations and the throat modulates voice production, subtle cardiovascular motions associated with these processes can be detected in human speech.
In video streams, when light hits the skin, subtle color changes from pulse-induced blood flow can be captured, as described by the Dichromatic Reflection Model (DRM) [50]. We define the Domain of Interest (DOI) of the facial areas as $\Pi\in\mathbb{R}^{N_{v}\times C\times H_{f}\times W_{f}}$, and $\Pi_{i,j}\in\Pi$ denotes the RGB pixels at the $i$-th row and the $j$-th column. To bridge the color with RGB values, we model the spectral relationship as:
@@ 3.1 Kinetics for Cardiac Learning
where $I(f)$ is the illumination spectral components, $\ast$ is the convolution operation, and $\Delta(f)$ is the reflection modulator, comprising specular reflection $\Delta_{s}(f)$ and diffuse reflection $\Delta_{d}(f)$. Specular reflection occurs at the epidermis level, while diffuse reflection penetrates into the hypodermis, reflecting off capillaries and blood vessels, encapsulating physiological spectrum $H(f)$. We further decompose $I(f)$ and $\Delta_{s}(f)$ into static and dynamic components, where dynamic components are denoted as $\mu(H(f),O(f))$ and $\nu(H(f),O(f))$, respectively. $O(f)$ is a set of irrelevant signals. $\mu(\cdot)$ and $\nu(\cdot)$ are transfer functions without analytic expressions. Our goal is to infer $h(t)$ from $\Pi$, where $h(t)$ is the temporal counterpart of the spectral representation $H(f)$.
@@
where $L(f)$ is the sound energy source. $R(f)$ is an acoustic filter creating formant, affected by the vocal tract’s physical attributes. Blood flow in surrounding vessels, particularly carotid arteries, influences the acoustic properties [67].These cardiovascular dynamics are encapsulated in the model by integrating the physiological signal $\hat{H}(f)$ into $R(f)$.
Observations: Existing video-based solutions [40, 21, 72, 41, 77, 61], though many, are trained on small datasets with controlled environments, e.g., PURE [51]. Their performances will degrade greatly when training and testing on more complicated datasets, e.g., MMPD [54]. As can be seen from Fig. 3, the existing video-based solutions cannot effectively capture the cardiac semantics across different body movements and light conditions. These results present a grand challenge for cardiac learning. Meanwhile, different light conditions and body movements will degrade the performance from the video-based approaches, where audio can help [67]. Therefore, our goal is to design a dedicated audio-visual network to extract those motions.
@@
3.2 Model Design
Given the underlying cardiac motions, we aim to devise a learning approach to extract $h(t)$. As shown in Fig. 4, the DOI pairs, i.e., frames $\Pi$ and audio clips $\Xi$, will be fed into video encoder $E_{v}$ and audio encoder $E_{a}$, respectively, to acquire the latent representation. Then we devise a fusion network to aggregate the two modalities. We will elaborate on the details of the design below.
3.2.1 Video Branch Design

### v2

@@ Abstract
Online Cardiac Monitoring (OCM) emerges as a compelling enhancement for the next-generation video streaming platforms. It enables various applications, including remote health, affective computing, and deepfake detection. Yet the physiological information encapsulated in the video streams has long been neglected. In this paper, we present the design and implementation of CardioLive, the first online cardiac monitoring system in video streaming platforms. We leverage the naturally co-existing video and audio streams and devise CardioNet, the first audio-visual network to learn the cardiac series. It incorporates multiple unique designs to extract temporal and spectral features, ensuring robust performance under realistic streaming conditions. To enable the Service-On-Demand OCM, we implement CardioLive as a plug-and-play middleware service and develop systematic solutions to practical issues including changing FPS and unsynchronized streams. Extensive evaluations demonstrate the effectiveness of our system. We achieve a Mean Squared Error of 1.79 BPM error, outperforming the video-only and audio-only solutions by 69.2% and 81.2%, respectively. CardioLive achieves average throughput of 115.97 and 98.16 FPS in Zoom and YouTube. We believe our work opens up new applications for video stream systems. Code is available at https://github.com/aiot-lab/CardioLive.
@@
I Introduction
Video streaming has exploded in recent years, with no slowdown in sight. From TikTok that have turned live video sharing into a global phenomenon, to Zoom, which has become synonymous with remote work, video streaming has woven itself into the fabric of our daily lives. The market is booming steadily [VideoStreamingSVoD], reflecting our collective appetite for real-time, interactive, and accessible content.
Online Cardiac Monitoring (OCM) can be one intriguing enhancement for the next-generation video streaming platforms. The rich tapestry of video and audio in streaming not only provides the context of actions, movement, and human activities, etc, but it also embeds subtle cardiac events, which have long been neglected in contemporary multimedia systems. Uncovering such physiological information would bring various benefits. For remote health, physicians could remotely access real-time cardiac data without the need for specialized equipment [HeartDiseaseSymptoms]. Similarly, in video gaming, displaying a player’s heart rate during live streams could add additional excitement and engagement for viewers [PulsoidRealtimeHearta]. OCM also plays a pivotal role in online conferences or interviews, where emotional responses inferred from cardiac data [wang2020you, sun2022estimating] could enrich interactions, making them more nuanced and meaningful. Furthermore, the potential for this technology extends into security and fraud detection against digital impersonation techniques like deepfakes [liu2019cardiocam, qi2020deeprhythm]. These applications of OCM underscore its potential to revolutionize video streaming, making it not just a tool for communication and entertainment but also a platform for health monitoring, affective computing, emotional intelligence, and security.
However, existing OCM either relies on specified hardware (e.g., heartbeat belt, wearables) or introduces additional modalities [liu2015tracking, yang2016monitoring, chen2021movi, xue2025ppg]. These approaches suffer from extra cost and are often misaligned with live streams. Moreover, sensing-based approaches necessitate active probing signals [wang2023df, qian2018acousticcardiogram, lyu2024ase], which are impractical in streaming. A video streaming system that seamlessly enables OCM in pervasive contexts without additional hardware still lacks.
@@ I Introduction
First, what information should we take from the video streaming system to monitor the cardiac activities? Existing works [chen2018deepphys, liu2020multi, niu2020video, li2023learning, yu2019remote, yu2019remoteCompress, yu2023physformer++, liu2023efficientphys, zou2024rhythmformer] on extracting heart rate from human faces focus on remote photoplethysmography (rPPG), which leverages solely video. These video-only solutions suffer from low illumination conditions, head movement, and orientation. Recent progress in cardiac vocal interfaces [xu2022hearing] inspires us to infer heart rate from human speech. However, audio signals are usually sensitive to noise interference and lack contextual background information, rendering them less robust in real-life scenarios and requiring user calibration. Conceptually, video provides detailed visual context while sound exhibits resilience to varying light conditions and body motions. Consequently, they offer complementary advantages to enhance cardiac monitoring. This motivates us to move beyond video-only or audio-only solutions and investigate new designs to combine the naturally co-existing video and audio streams.
@@
Second, how to tackle real-world problems to make this system robust and accurate? Unveiling the cardiac activity from video and audio is challenging. The information is easy to be overshadowed by more prominent body movements, environmental dynamics, and/or ambient noise. Previous works [liu2020multi, niu2020video, li2023learning, yu2023physformer++, liu2023efficientphys, zou2024rhythmformer] primarily evaluate models on well-controlled datasets featuring static subjects under optimized light conditions and viewing angles, which simplifies the problems yet becomes unrealistic in real-world settings. The task gets even more challenging when deployed in live video streaming environments, due to the discrepancies in frame rates and degraded image quality. To deliver an accurate and robust system in practice, novel techniques are desired to effectively discern subtle cardiac signals amidst various disturbances while combating fluctuating frame rates and drifted misalignment of the streams.
@@
Third, how to enable Service-On-Demand (SoD) cardiac monitoring in video streaming system Despite the promise of the integration, enabling SoD for users poses significant challenges due to the complexity of modern video streaming systems. These platforms vary widely, encompassing formats such as conferences [SkypeStayConnected, TeamsChannelsMicrosoft, OnePlatformConnect], Video-On-Demand (VoD) [StreamTVMovies, NetflixSingaporeWatch], live streaming [YouTube, ExploreFindYour], etc, each with its own technical and operational nuances. These providers balance the demands of real-time data processing with the need for immediate accessibility and minimal latency while not interfering with the original streams. At the same time, deploying our service on edge (e.g., browsers) benefits from preserving privacy, while getting access to the data yields another challenge. One naive way is to deploy our models over the WebRTC peers, but it lacks scalability and versatility. To this end, we are motivated to establish a plug-and-play service that can be seamlessly integrated into video streaming systems, whether hosted on servers or edges.
@@
In this paper, we present CardioLive, the first-of-its-kind OCM system, that can continuously infer the heart rate in video streaming systems. At the core of CardioLive, we design a novel audio-video network, CardioNet, that effectively learns the nuanced cardiac activities from facial regions and human voices. We further devise systematic solutions to deploy CardioLive as a middleware service to support the SoD online cardiac monitoring. We introduce practical techniques to handle changing FPS and unsynchronized streams. Through in-depth analyses of the streaming architectures, we design effective data hooks and a novel packet buffer, which can be easily integrated with various video streaming systems.
@@
Extensive experiments have been done to validate the effectiveness of CardioLive. Our evaluation results show that CardioLive achieves a mean absolute error (MAE) of 1.79 BPM and root mean square error (RMSE) of 3.25 BPM, largely outperforming the video-only solutions by 69.2% in MAE and 61.4% in RMSE, and the audio-only solution by 81.2% in MAE and 76.8% in RMSE. As for CardioLive service, we implement our system on two ends, a meeting platform (Zoom) and a content provider (YouTube), respectively. We achieve the overall throughput of 115.97 FPS and 98.16 FPS for each platform, respectively, ensuring smooth updates without disrupting the original streams. These results highlight the robustness and accuracy of CardioLive, confirming its potential for widespread application in video streaming systems.
@@
❸ We implement CardioLive as a service-based plug-and-play middleware that can seamlessly be integrated into mainstream platforms for real-time streaming.
II Design Scope
@@
Accessibility: In many video streaming scenarios, such as live product demonstrations on TikTok or Zoom interviews, using wearables or additional hardware is often impractical. OCM can overcome this problem by leveraging modalities that already exist within video streams, thereby increasing accessibility for audiences and facilitating broader engagement. It also promises wider dissemination of remote health, offering device-free cardiac monitoring compared to the latest work [chen2024exploring] that relies on earphones.
@@
Enhanced Analytical Abilities: While there exist alternative approaches for tasks including affective computing [mottelson2016affect, prajwal2023towards, wu2020emo, ahmad2024detecting] and deepfake detection [yang2023avoid, demir2021deep], the cardiac signal shows a strong correlation with them [wascher2021heart, prajwal2023towards], by capturing the subtle changes in heart rate. In this context, OCM provides an additional verification layer in a real-time and continuous manner, allowing experts to analyze behaviors. This analysis can help determine if someone is lying, happy, nervous, or engaging in deceptive behavior.
@@
Entertainment: Our work also presents a distinct chance for augmented entertainment. With the rise of live streaming, the audience can access the heart rates of celebrities, which opens up a new world for existing viewing experiences.
Figure 2: Kinetics of Cardiac Learning.
@@
Despite the potential, there are no existing solutions capable of achieving this integration without additional hardware. In this work, we focus on addressing this gap by leveraging the co-existence of audio and video signals, specifically in scenarios where a speaker is talking. This can be common in both entertainment and telehealth use cases. At the core of OCM is the accurate prediction of cardiac information. Our system should robustly detect the heart rate from the video streaming systems by hooking the video and audio chunks. Once cardiac data is acquired, it can be further analyzed for various downstream tasks, including affective computing, remote health monitoring, and deepfake detection. Yet how cardiac monitoring is used for downstream tasks (e.g., emotions, lies, etc) is not the focus of this paper.
Audio-Video Pair: We intend to integrate the video and audio information for cardiac monitoring. Leveraging the natural co-existence of audio and video modalities offers contemporary benefits as follows: ❶ Ubiquity: Video and audio streams are the most fundamental components in video streaming systems, while no additional hardware is needed. ❷ Feasibility: Both video and audio data contain the cardiac information (discussed in §III-A). ❸ Complementarity: Audio and video offer different strengths and weaknesses. Audio is less interfered with by motion and light but is sensitive to noises. Video is more robust to noise but will fail in various body movements and non-optimized view angles. We will elaborate on the detailed analyses in §III. We argue that in our primary target application scenarios—such as video conferencing, live streaming, and remote healthcare—human speech is inherently present alongside video. Our goal is to fully leverage the potential of these naturally coexisting signals. Additionally, our system is well-designed to seamlessly fall back to a video-only solution when audio quality degrades.
@@
CardioLive as a service: To deploy such an OCM system, a straightforward way is to build a self-hosted WebRTC service, which, however, does not scale to existing streaming systems. Therefore, for the sake of versatility, we establish a microservice to host CardioLive for seamless integration with mainstream video streaming platforms.
@@
In a nutshell, the audio-video pair appears to be an attractive choice for ubiquitous and practical OCM, yet it entails numerous challenges to build an accurate and robust multi-modal algorithm and system. We will present our model design in §III and leave the system implementation in §IV.
@@
Figure 4: The performances of video-based approaches vary under different body movements and light conditions.
III Model Design
III-A Kinetics for Cardiac Learning
Principles: In this section, we introduce the principles of extracting cardiac information from video and audio data, as shown in Fig. 2. The fundamental concept revolves around the variations in blood pressure caused by cardiac activities, which manifest as quasi-periodic deformations of blood vessels. Since blood vessels circulate blood throughout the body, including the face, lungs, and throat, we can infer heart activity in these areas through video and audio analysis. Specifically, when a light illuminates the skin, subtle color variations caused by pulse-induced blood flow can be captured through video streams. Additionally, as the lungs supply airflow for vocal fold vibrations and the throat modulates voice production, subtle cardiovascular motions associated with these processes can be detected in human speech.
In video streams, when light hits the skin, subtle color changes from pulse-induced blood flow can be captured, as described by the Dichromatic Reflection Model (DRM) [shafer1985using]. We define the Domain of Interest (DOI) of the facial areas as $\Pi\in\mathbb{R}^{N_{v}\times C\times H_{f}\times W_{f}}$, and $\Pi_{i,j}\in\Pi$ denotes the RGB pixels at the $i$-th row and the $j$-th column. To bridge the color with RGB values, we model the spectral relationship as:
@@ III-A Kinetics for Cardiac Learning
where $I(f)$ is the illumination spectral components, $\ast$ is the convolution operation, and $\Delta(f)$ is the reflection modulator, comprising specular reflection $\Delta_{s}(f)$ and diffuse reflection $\Delta_{d}(f)$. Specular reflection occurs at the epidermis level, while diffuse reflection penetrates into the hypodermis, reflecting off capillaries and blood vessels, encapsulating the physiological spectrum $H(f)$. We further decompose $I(f)$ and $\Delta_{s}(f)$ into static and dynamic components, where dynamic components are denoted as $\mu(H(f),O(f))$ and $\nu(H(f),O(f))$, respectively. $O(f)$ is a set of irrelevant signals. $\mu(\cdot)$ and $\nu(\cdot)$ are transfer functions without analytic expressions. Our goal is to infer $h(t)$ from $\Pi$, where $h(t)$ is the temporal counterpart of the spectral representation $H(f)$.
@@
where $L(f)$ is the sound energy source. $R(f)$ is an acoustic filter creating formant, affected by the vocal tract’s physical attributes. Blood flow in surrounding vessels, particularly carotid arteries, influences the acoustic properties [xu2022hearing]. These cardiovascular dynamics are encapsulated in the model by integrating the physiological signal $\hat{H}(f)$ into $R(f)$.
Observations: Existing video-based solutions [liu2020multi, chen2018deepphys, yu2019remote, liu2023efficientphys, zou2024rhythmformer, wang2016algorithmic], though many, are trained on small datasets with controlled environments, e.g., PURE [stricker2014non]. Their performances will degrade greatly when training and testing on more complicated datasets, e.g., MMPD [10340857]. As can be seen from Fig. 4, the existing video-based solutions cannot effectively capture the cardiac semantics across different body movements and light conditions. These results present a grand challenge for cardiac learning. Meanwhile, different light conditions and body movements will degrade the performance from the video-based approaches, where audio can help [xu2022hearing]. Therefore, our goal is to design a dedicated audio-visual network to extract those motions.
@@
III-B CardioNet Design
As shown in Fig. 5, the DOI pairs, i.e., frames $\Pi$ and audio clips $\Xi$, will be fed into video encoder $E_{v}$ and audio encoder $E_{a}$, respectively, followed by a fusion network to aggregate the two modalities.
III-B1 Video Branch Design